# Enhancing Time Series Forecasting via Multi-Level Text Alignment with LLMs

### v1

@@ Abstract
The adaptation of large language models (LLMs) to time series forecasting poses unique challenges, as time series data is continuous in nature, while LLMs operate on discrete tokens. Despite the success of LLMs in natural language processing (NLP) and other structured domains, aligning time series data with language-based representations while maintaining both predictive accuracy and interpretability remains a significant hurdle. Existing methods have attempted to reprogram time series data into text-based forms, but these often fall short in delivering meaningful, interpretable results. In this paper, we propose a multi-level text alignment framework for time series forecasting using LLMs that not only improves prediction accuracy but also enhances the interpretability of time series representations. Our method decomposes time series into trend, seasonal, and residual components, which are then reprogrammed into component-specific text representations. We introduce a multi-level alignment mechanism, where component-specific embeddings are aligned with pre-trained word tokens, enabling more interpretable forecasts. Experiments on multiple datasets demonstrate that our method outperforms state-of-the-art models in accuracy while providing good interpretability.
@@ 1 Introduction
Time series forecasting, which involves predicting future values based on historical data, has numerous practical applications, such as demand planning, inventory optimization, energy load forecasting, and climate modeling [7, 13, 14, 6]. Traditionally, these tasks demand substantial domain expertise and carefully designed models tailored to specific datasets. However, recent advancements in pre-trained large language models (LLMs), such as GPT-4 [1] and LLaMA [23], have achieved remarkable success in natural language processing (NLP) and demonstrated potential in handling complex, structured domains. This raises a compelling question: how can these powerful pre-trained LLMs be effectively adapted for time series forecasting?
LLMs, trained on vast and diverse text corpora, provide a powerful foundation for various downstream tasks, requiring only minimal task-specific prompt engineering or fine-tuning. This flexibility has sparked a growing interest in leveraging LLMs for time series analysis. For example, methods like Promptcast [28] and LLMTime [8] reformulate numerical inputs and outputs into prompts, treating time series forecasting as a sentence-to-sentence task, which enables the direct application of LLMs. Meanwhile, approaches like TEMPO [2] and GPT4TS [33] take a different route by fine-tuning pre-trained LLMs, modifying components such as the Add&Norm layers and positional embeddings, further demonstrating LLMs’ adaptability for time series forecasting.
@@
Figure 1: Cross-modality time series embeddings of $(a)$ contrast learning of text-prototype-aligned time series embeddings, $(b)$ text prototypes reprogramming, $(c)$semantic informed prompt and and $(d)$ proposed anchors alignment of our multi-level alignment.
In this paper, we address the challenge of interpretability in LLM-based time series forecasting by developing an interpretable multi-level text alignment framework while preserving the backbone model. Our approach consists of two key principles for effective time series representation learning: (a) modeling specific time series components such as trend, seasonality, and residuals, and (b) deriving interpretable explanations from the inherent properties of time series data through multi-level text alignment. Specifically, we decompose the time series input into three additive components—trend, seasonality, and residuals—using locally weighted scatterplot smoothing (LOESS) [5]. These components are then reprogrammed into component-specific anchors that better align with the language capabilities of LLMs in Figure 1(d). Additionally, we employ component-specific prompts to guide the generation of learnable continuous vector representations that encode temporal knowledge of each component.
In summary, the main contributions of this paper are as follows: (1) We propose an interpretable multi-level text alignment framework for time series forecasting using LLMs, while keeping the backbone model unchanged. (2) Our method leverages this multi-level alignment to map decomposed time series components—trend, seasonality, and residuals—into distinctive, informative joint representations. The aligned trend-specific anchors enhance the interpretability of LLMs, while the aligned seasonality and residual prototypes improve the overall representation of the input time series. (3) Experimental results on multiple datasets validate the superiority of our model over state-of-the-art approaches, highlighting the effectiveness of interpretable multi-level text alignment.
@@ 2 Related work
2.1 Pre-trained Large Language Models for Time series.
The recent advancements in Large Language Models (LLMs) have opened up new opportunities for time series modeling. LLMs like T5 [21], GPT-2 [20], GPT-4 [1], and LLaMA [23] have demonstrated impressive capabilities in understanding complex dependencies in heterogeneous textual data and generating meaningful outputs. Recently, there has been growing interest in exploring how to transfer the knowledge embedded in these pre-trained LLMs to the time series domain [11, 9]. For instance, [28] converts time series data into text sequences, achieving promising results. Other works, such as [33, 8], tokenize time series data into overlapping patches and strategically fine-tune LLMs for time series forecasting tasks. Similarly, recent works such as [2, 19] decompose time series data and use retrieval-based prompts to enhance fine-tuning of pre-trained LLMs. However, these approaches often fall short of delivering interpretable results and tend to treat time series as mere sequences of tokens, overlooking their inherent temporal structures. Converting numerical data to text without sufficient alignment to temporal dynamics can lead to inaccurate predictions and a lack of transparency in the model’s decision-making process, especially for multivariate time series. Our work introduces an interpretable multi-level text alignment framework to align time series components with anchors while keeping pre-trained LLMs intact.
2.2 Time Series Aligned Embeddings
[…]
@@ 3 Methodology
Our approach focuses on enhancing the interpretability of large language models (LLMs) for time series data through multi-level aligned embeddings. As illustrated in Figure 2, the proposed framework consists of four core modules: (1) time series input decomposition, (2) multi-level text alignment, (3) component-specific prompts, and (4) output projection. The process begins by partitioning a multivariate time series into $N$ univariate time series, each processed independently. The $i$-th series, denoted as $\mathbf{X}^{(i)}\in\mathbb{R}^{1\times L}$, undergoes a series of steps including decomposition, normalization, patching, and embedding before being aligned with anchor points and text prototypes. To enhance the LLM’s reasoning capability on time series data, we introduce component-specific prompts along with the aligned embeddings, enabling the model to generate meaningful output representations. These representations are then projected through an output linear layer to produce the final forecasts, $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$. With the primary objective of improving interpretability, we utilize GPT-2 [20], employing its first six layers as the backbone model for time series forecasting without fine-tuning the foundational model.
@@ 3.1 Problem statement
Given the observed values over the previous $L$ timestamps, the task of multivariate time series forecasting is to predict the values for the next $H$ timestamps. Formally, this can be represented as:
@@
where $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$ is the vector of $H$-step prediction from timestamp $t$ of channel $i$ corresponding to the $i$-th feature. Given the historical values $\mathbf{x}_{t-L}^{(i)},\dots,\\ \mathbf{x}_{t-1}^{(i)}$, a large language model $F$ uses prompt $\mathbf{V}^{(i)}$ to make these predictions. Leveraging the strong reasoning capabilities of pre-trained large language models, we aim to align time series data with text to enable LLMs to interpret the input series and accurately forecast the $H$ future steps, with the overall objective of minimizing the mean square errors between the ground truths and predictions, expressed as:
@@ 3.2 Time series input decomposition
where $i$ refers to the feature index for multivariate time series input. The trend component $X_{T}\in\mathbb{R}^{N\times L}$, capturing the underlying long-term patterns in the data, is expressed as $X_{T}=\frac{1}{m}\sum_{j=-k}^{k}X_{t+j}$, where $m=2k+1$ and $k$ is the averaging step size. The seasonal component $X_{S}\in\mathbb{R}^{N\times L}$ reflects the repeating short-term cycles and can be estimated after removing the trend. The residual component $X_{R}\in\mathbb{R}^{N\times L}$ represents the remainder of the data once the trend and seasonal elements have been extracted. There are multiple methods available for performing additive seasonal-trend decomposition. One common approach is the classical additive seasonal-trend decomposition, which first extracts the long-term trend using moving averages. The seasonal component is then estimated by averaging the detrended time series, and the residual is obtained by subtracting the estimated trend and seasonal components. Another widely used method is the Seasonal-Trend decomposition using Loess (STL) [5]. The choice of decomposition method in this paper is determined based on validation results.
Following the approach outlined in [17], we patch the decomposed components of the time series. Specifically, for the $i$-th normalized trend component, we obtain the patched token $\textbf{P}^{(i)}_{T}\in\mathbb{R}^{L_{P}\times K}$, where $L_{P}$ represents the patch length and $K=\lfloor\frac{(L-L_{P})}{s}\rfloor+2$ denotes the number of patches, with $s$ is the stride. Similarly, we apply this patching process to the seasonal and residual components, obtaining patched tokens $\textbf{P}^{(i)}_{S}$ and $\textbf{P}^{(i)}_{R}$, respectively. These patched tokens are then fed into the multi-level text alignment module to produce aligned time series embeddings.
@@ 3.3 Multi-level text alignment
Here we reprogram patch embeddings into the LLMs’ pre-training data representation space to align the modalities of time series and natural language to activate the backbone’s time series understanding and reasoning capabilities. Naively, we can align the token embedding of time series and text using similarity estimation. Although time series tokens lack text annotation, we can place their embedding near typical text descriptions of time series. Thus, it is intuitively expected that various time series tokens can represent various descriptive words such as up, down, stable, and so on. However, the pre-trained word token embedding space is vast and dense, and the selection of text prototypes (patterns) is often highly relaxed, sometimes even involving random words unrelated to time series or clusters of pre-trained word tokens [22, 19, 10], which leads to poor interpretability.
@@
As illustrated in the top-right of Figure 2, our multi-level text alignment aims to give a connection between anchors and trend patches. The selected anchors are sparse. We reprogram seasonal and residual with text prototypes connecting time series patches with a more dense reprogramming space. To realize this, we employ a multi-head cross-attention layer for each component. Specifically, for $i$-th input feature, we define query matrices $\mathbf{Q}_{T}^{(i)}=\textbf{P}_{T}^{(i)}\textbf{W}^{Q(i)}_{T}$, key matrices $\mathbf{K}_{T}^{(i)}=\mathbb{W}_{trend}\textbf{W}^{K(i)}_{T}$, value matrices $\mathbf{V}_{T}^{(i)}=\mathbb{W}_{trend}\textbf{W}^{V(i)}_{T}$ for trend; query matrices $\mathbf{Q}_{S}^{(i)}=\textbf{P}_{S}\textbf{W}^{Q(i)}_{S}$, key matrices $\mathbf{K}_{S}^{(i)}=\textbf{E}^{{}^{\prime}}_{seasonal}\textbf{W}^{K(i)}_{S}$, value matrices $\mathbf{V}_{S}^{(i)}=\textbf{E}^{{}^{\prime}}_{seasonal}\textbf{W}^{V(i)}_{S}$ for seasonal; query matrices $\mathbf{Q}_{R}^{(i)}=\textbf{P}_{R}\textbf{W}^{Q(i)}_{R}$, key matrices $\mathbf{K}_{R}^{(i)}=\textbf{E}^{{}^{\prime}}_{residual}\textbf{W}^{K(i)}_{R}$, value matrices $\mathbf{V}_{R}^{(i)}=\textbf{E}^{{}^{\prime}}_{residual}\textbf{W}^{V(i)}_{R}$ for residual. Through multi-head attention, we reprogram each time series component. For example, the trend after multi-head attention is defined as:
$\mathbf{Z}_{T}^{(i)}=\text{ATTENTION}(\mathbf{Q}_{T}^{(i)},\mathbf{K}_{T}^{(i)},\mathbf{V}_{T}^{(i)})=\text{SOFTMAX}\left(\frac{\mathbf{Q}_{T}^{(i)}\mathbf{K}_{T}^{(i)\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{T}^{(i)},$ (4)
where $d_{k}$ is the dimension of each head in the multi-head attention module. After the multi-head attention step, each component is linearly projected to align the hidden dimensions with the backbone model.
3.4 Component-specific prompts
Prompting techniques have proven highly effective across various tasks by leveraging task-specific knowledge encoded in prompts. This success stems from prompts providing a structured framework that aligns the model’s output with desired objectives, improving accuracy and coherence. However, directly translating time series into natural language poses challenges, complicating the creation of instruction-following datasets and effective on-the-fly prompting [27]. Recent advances show that prompts can enrich input context and guide the transformation of reprogrammed time series patches [10]. To leverage the semantic information of time series components, we propose a component-specific prefix prompting strategy. This includes three elements: dataset context, input statistical features, and component-specific task instructions for trend, seasonal, and residual components. For instance, the task description ’forecast the next 96 steps given the previous 512 steps [trend, seasonal, residual]’ serves as a template for our task instructions, which are then concatenated with the corresponding component data.
3.5 Output projection
After packing and forwarding the component-specific prompts and embeddings through the frozen backbone LLM, we retain the embedding for each component and apply a linear projection to the output representation. By denormalizing and summing these representations, we derive the final forecasts $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$.
@@ 4 Experiments
In our experiments, the proposed method outperforms state-of-the-art forecasting approaches across various benchmarks, including long-term, short-term, and few-shot forecasting. For a fair comparison, we follow the configurations outlined in [25] across all baselines, utilizing a unified evaluation pipeline [^1] https://github.com/thuml/Time-Series-Library. Our code will be made available on GitHub upon the acceptance of the paper.
Baselines. We compare with the SOTA time series models and cite their performance from [33, 4] if applicable. The SOTA includes a set of Transformer-based methods, i.e., PatchTST [17], ETSformer [24], Non-Stationary Transformer [16], FEDformer [32], Autoformer [26], Informer [31], and Reformer [12]. We also select a set of non-transformer-based techniques, i.e., DLinear [29], TimesNet [25], N-BEATS [18], and LightTS [30]. Finally, four methods are based on LLMs, i.e., TimeLLM [10], LLM4TS [4], GPT4TS [33], and LLMTime [8]. Aligned with the GPT4TS configuration [33], we utilize only the first 6 layers of the 12-layer GPT-2 base [20] as the backbone model of ours and TimeLLM.
4.1 Long-term Forecasting
Setup. For long-term forecasting, we evaluate on ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity(ECL), and Traffic, which have been widely adopted as benchmarking datasets for long-term forecasting works [25]. The input time series length $L$ is set as 512, and we evaluate across four prediction horizons: $H\in\{96,192,336,720\}$. The evaluation metrics include mean square error (MSE) and mean absolute error (MAE).
Results. Table 1 presents the performance of various time series forecasting models on MSE and MAE metrics across different prediction horizons on multiple benchmarks. Our proposed model consistently outperforms existing baselines, demonstrating superior performance on average across most datasets and prediction lengths. This highlights the applicability of multi-level text alignment. Notably, our comparison with TimeLLM—a recent work leveraging text prototype reprogramming to align time series with text tokens—is significant. Specifically, our model achieves substantial improvements on the Weather and ETTm1 datasets, exceeding the best-performing LLM-based model, LLM4TS, by 23.3% and 26.8%, respectively, in terms of MSE. Additionally, it records the lowest error rates across numerous dataset-prediction length configurations. These results suggest that integrating LLMs with multi-level text alignment can significantly enhance the accuracy of long-term time series forecasting.
Table 1: Long-term forecasting results for {96, 192, 336, 720} horizons. Lower values indicate better performance. bold: the best, underline: second best.
@@ 4.1 Long-term Forecasting
Datasets $\setminus$ Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.355 0.404 0.384 0.407 0.371 0.394 0.376 0.397 0.375 0.399 0.370 0.399
192 0.426 0.445 0.423 0.434 0.403 0.412 0.416 0.418 0.405 0.416 0.413 0.421
336 0.434 0.449 0.435 0.447 0.420 0.422 0.442 0.433 0.439 0.443 0.422 0.436
720 0.480 0.493 0.439 0.463 0.422 0.444 0.477 0.456 0.472 0.490 0.447 0.466
Avg. 0.424 0.448 0.420 0.438 0.404 0.418 0.428 0.426 0.422 0.437 0.413 0.435
ETTh2 96 0.260 0.336 0.295 0.355 0.269 0.332 0.285 0.342 0.289 0.353 0.274 0.336
192 0.333 0.375 0.376 0.410 0.328 0.377 0.354 0.389 0.383 0.418 0.339 0.379
336 0.369 0.408 0.376 0.412 0.353 0.396 0.373 0.407 0.448 0.465 0.329 0.380
720 0.444 0.455 0.410 0.442 0.383 0.425 0.406 0.441 0.605 0.551 0.379 0.422
Avg. 0.378 0.408 0.364 0.403 0.333 0.383 0.355 0.394 0.431 0.446 0.330 0.379
ETTm1 96 0.117 0.232 0.297 0.349 0.285 0.343 0.292 0.346 0.299 0.343 0.290 0.342
192 0.198 0.298 0.336 0.373 0.324 0.366 0.332 0.372 0.335 0.365 0.332 0.369
336 0.301 0.360 0.362 0.390 0.353 0.385 0.366 0.394 0.369 0.386 0.366 0.392
720 0.389 0.411 0.410 0.421 0.408 0.419 0.417 0.421 0.425 0.421 0.416 0.425
Avg. 0.251 0.325 0.351 0.383 0.343 0.378 0.352 0.383 0.357 0.378 0.351 0.380
ETTm2 96 0.095 0.200 0.177 0.264 0.165 0.254 0.173 0.262 0.167 0.269 0.165 0.255
192 0.174 0.263 0.253 0.312 0.220 0.292 0.229 0.301 0.224 0.303 0.220 0.292
336 0.243 0.313 0.285 0.345 0.268 0.326 0.286 0.341 0.281 0.342 0.274 0.329
720 0.343 0.380 0.366 0.390 0.350 0.380 0.378 0.401 0.297 0.421 0.362 0.385
Avg. 0.214 0.289 0.270 0.328 0.251 0.313 0.267 0.326 0.267 0.333 0.255 0.315
Weather 96 0.059 0.125 0.158 0.210 0.147 0.196 0.162 0.212 0.176 0.237 0.149 0.198
192 0.115 0.188 0.191 0.240 0.191 0.238 0.204 0.248 0.220 0.282 0.194 0.241
336 0.211 0.263 0.247 0.284 0.241 0.277 0.254 0.286 0.265 0.319 0.245 0.282
720 0.299 0.327 0.319 0.334 0.313 0.329 0.326 0.337 0.333 0.362 0.314 0.334
Avg. 0.171 0.226 0.229 0.267 0.223 0.260 0.237 0.271 0.248 0.300 0.225 0.264
ECL 96 0.116 0.221 0.137 0.237 0.128 0.223 0.139 0.238 0.140 0.237 0.129 0.222
192 0.145 0.250 0.150 0.249 0.146 0.240 0.153 0.251 0.153 0.249 0.150 0.240
336 0.167 0.271 0.168 0.266 0.163 0.258 0.169 0.266 0.169 0.267 0.163 0.259
720 0.209 0.307 0.203 0.293 0.200 0.292 0.206 0.297 0.203 0.301 0.197 0.290
Avg. 0.159 0.262 0.164 0.261 0.159 0.253 0.167 0.263 0.166 0.263 0.161 0.252
Traffic 96 0.255 0.229 0.380 0.277 0.372 0.259 0.388 0.282 0.410 0.282 0.360 0.249
192 0.332 0.258 0.399 0.288 0.391 0.265 0.407 0.290 0.423 0.287 0.379 0.256
336 0.370 0.273 0.408 0.290 0.405 0.275 0.412 0.294 0.436 0.296 0.392 0.264
720 0.428 0.301 0.445 0.308 0.437 0.292 0.450 0.312 0.466 0.315 0.432 0.286
Avg. 0.346 0.265 0.408 0.290 0.401 0.273 0.414 0.294 0.433 0.295 0.390 0.263
4.2 few-shot forecasting
Setups. LLMs have recently shown impressive few-shot learning capabilities [15]. To evaluate performance in the few-shot forecasting setting, we follow the experimental setup outlined in [33], allowing us to assess whether the model can generate accurate forecasts with limited training data. In these experiments, we use only the first 10% of the training data.
Results. The brief 10% few-shot learning results in Table 2 and full results in Table 5 demonstrate that our model significantly outperforms all baseline methods across most datasets. We attribute this success to the effective knowledge activation achieved through multi-level text alignment. Specifically, our model improves few-shot learning performance on the Weather and Traffic datasets by 11.9% and 5%. When trained with only 10% of the data, LLM-based methods substantially outperform other baselines, which are trained from scratch and thus limited by the smaller training set. In contrast, LLM-based models can leverage pre-trained knowledge and align it with time series embeddings to enhance representation.
Table 2: Few-shot learning on 10% training data. All results are averaged from four different forecasting horizons: $H\in\{96,192,336,720\}$. Lower values indicate better performance.
Methods Ours TimeLLM LLM4TS GPT4TS DLinear PatchTST

### v2

@@ Abstract
The adaptation of large language models (LLMs) to time series forecasting poses unique challenges, as time series data is continuous in nature, while LLMs operate on discrete tokens. Despite the success of LLMs in natural language processing (NLP) and other structured domains, aligning time series data with language-based representations while maintaining both predictive accuracy and interpretability remains a significant hurdle. Existing methods have attempted to reprogram time series data into text-based forms, but these often fall short in delivering meaningful, interpretable results. In this paper, we propose a multi-level text alignment framework for time series forecasting using LLMs that not only improves prediction accuracy but also enhances the interpretability of time series representations. Our method decomposes time series into trend, seasonal, and residual components, which are then reprogrammed into component-specific text representations. We introduce a multi-level alignment mechanism, where component-specific embeddings are aligned with pre-trained word tokens, enabling more interpretable forecasts. Experiments on multiple datasets demonstrate that our method outperforms state-of-the-art models in accuracy while providing good interpretability. Our code is available at https://github.com/ztb-35/MLTA.
@@ 1 Introduction
Large language models (LLMs), trained on vast and diverse text corpora, provide a powerful foundation for various downstream tasks, requiring only minimal task-specific prompt engineering or fine-tuning. This flexibility has sparked a growing interest in leveraging LLMs for time series analysis. For example, methods like Promptcast [14] and LLMTime [5] reformulate numerical inputs and outputs into prompts, treating time series forecasting as a sentence-to-sentence task, which enables the direct application of LLMs. Meanwhile, approaches like TEMPO [2] and GPT4TS [16] take a different route by fine-tuning pre-trained LLMs, modifying components such as the Add&Norm layers and positional embeddings, further demonstrating LLMs’ adaptability for time series forecasting.
@@
[…]
In this paper, we address the challenge of interpretability in LLM-based time series forecasting by developing an interpretable multi-level text alignment framework while preserving the backbone model. Our approach consists of two key principles: (a) modeling time series components—trend, seasonality, and residuals, and (b) deriving interpretable explanations via multi-level text alignment. Using additive decomposition methods, like LOESS [4], we decompose time series into three additive components, reprogramming them into component-specific anchors for better LLM alignment (Figure 1(d)). Additionally, we employ component-specific prompts to guide the generation of learnable continuous vector representations that encode temporal knowledge of each component.
In summary, the main contributions of this paper are as follows: (1) We propose an interpretable multi-level text alignment framework for time series forecasting using LLMs while keeping the backbone model unchanged. (2) Our method leverages this multi-level alignment to map decomposed time series components—trend, seasonality, and residuals—into distinctive, informative joint representations. The aligned trend-specific anchors enhance the interpretability of LLMs, while the aligned seasonality and residual prototypes improve the overall representation of the input time series. (3) Experimental results on multiple datasets validate the superiority of our model over state-of-the-art approaches, highlighting the effectiveness of interpretable multi-level text alignment.
@@ 2 Related work
Recent advancements in LLMs like GPT-2 [9] and LLaMA [11] have enabled new possibilities for time series modeling. Efforts to adapt LLMs for this domain include converting time series into text [14], tokenizing data into patches for fine-tuning [16, 5], and leveraging retrieval-based prompts [2, 8]. However, these approaches often lack interpretability, treating time series as mere token sequences and ignoring temporal structures. This misalignment can lead to inaccurate predictions and reduced model transparency, particularly for multivariate time series. To address this, we propose an interpretable multi-level text alignment framework that preserves the integrity of pre-trained LLMs while effectively aligning time series components with anchors.
A key challenge in adapting LLMs for time series forecasting is aligning continuous data with discrete token embeddings. Inspired by prototype-level contrast methods, works like [10] select text embeddings as prototypes (Figure 1(a)), while [6] reprogram time series using source modality prompts without altering the backbone LLM (Figure 1(b)). These methods follow a "time series → pattern → text" paradigm, but their arbitrary prototype selection may not capture the true characteristics of the data [10]. More detailed related works see full paper [1]. To overcome this, our approach uses time series-specific anchors (e.g., "increase, decrease, stable") to guide token embedding learning (Figure 1(d)), enhancing both interpretability and forecasting performance.
@@ 3 Methodology
Our approach enhances the interpretability of LLMs for time series data using multi-level aligned embeddings. As shown in Figure 2, the framework consists of four core modules: (1) time series decomposition, (2) multi-level text alignment, (3) component-specific prompts, and (4) output projection. A multivariate time series is first partitioned into $N$ univariate series, each processed independently as $\mathbf{X}^{(i)}\in\mathbb{R}^{1\times L}$. The series undergoes decomposition, normalization, patching, and embedding before alignment with anchor points and text prototypes. To enhance reasoning, we introduce component-specific prompts, enabling the model to generate meaningful representations, which are projected through a linear layer to produce the final forecasts, $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$. We utilize GPT-2 [9], employing its first six layers as the backbone for forecasting without fine-tuning.
@@ 3.1 Problem statement
Given the previous $L$ observed values, the multivariate time series forecasting is predicts the values for the next $H$ timestamps. It can be formally as:
@@
where $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$ is the vector of $H$-step prediction from timestamp $t$ of channel $i$ corresponding to the $i$-th feature. Given the historical values $\mathbf{x}_{t-L}^{(i)},\dots,\\ \mathbf{x}_{t-1}^{(i)}$, a large language model $F$ uses prompt $\mathbf{V}^{(i)}$ to make these predictions. Leveraging the strong reasoning capabilities of pre-trained LLMs, we aim to align time series data with text to enable LLMs to interpret the input series and accurately forecast the $H$ future steps, with the overall objective expressed as:
@@ 3.2 Time series input decomposition
where $i$ refers to the feature index for multivariate time series input. The trend component $X_{T}\in\mathbb{R}^{N\times L}$, capturing the underlying long-term patterns in the data, is expressed as $X_{T}=\frac{1}{m}\sum_{j=-k}^{k}X_{t+j}$, where $m=2k+1$ and $k$ is the averaging step size. The seasonal component $X_{S}\in\mathbb{R}^{N\times L}$ reflects the repeating short-term cycles and can be estimated after removing the trend. The residual component $X_{R}\in\mathbb{R}^{N\times L}$ represents the remainder. Additive decomposition can be performed using various methods. A common approach first extracts the long-term trend via moving averages, estimates the seasonal component by averaging the detrended series, and obtains the residual by subtraction. Another widely used method is Seasonal-Trend decomposition using Loess (STL) [4]. The decomposition method in this paper is chosen based on validation results.
Following the approach outlined in [7], we patch the decomposed components of the time series. Specifically, for the $i$-th normalized trend component, we obtain the patched token $\textbf{P}^{(i)}_{T}\in\mathbb{R}^{L_{P}\times K}$, where $L_{P}$ represents the patch length and $K=\lfloor\frac{(L-L_{P})}{s}\rfloor+2$ denotes the number of patches, with $s$ is the stride. Similarly, we apply this patching process to the seasonal and residual components, obtaining patched tokens $\textbf{P}^{(i)}_{S}$ and $\textbf{P}^{(i)}_{R}$, respectively. These patches are then fed into the multi-level text alignment to produce aligned time series embeddings.
@@ 3.3 Multi-level text alignment
We reprogram patch embeddings into the LLMs’ pre-training space to align time series and natural language modalities, enhancing the backbone’s reasoning capabilities. A naive approach aligns time series and text embeddings via similarity estimation, placing time series tokens near relevant text descriptors (e.g., up, down, stable). However, the pre-trained word embedding space is vast and dense, and selecting appropriate text prototypes is often arbitrary, sometimes involving unrelated words or clusters [10, 8, 6], leading to poor interpretability.
@@
As illustrated in the top-right of Figure 2, our multi-level text alignment aims to give a connection between anchors and trend patches. The selected anchors are sparse. We reprogram seasonal and residual with text prototypes connecting time series patches with a more dense reprogramming space. To realize this, we employ a multi-head cross-attention layer for each component. Specifically, for $i$-th input feature, we define query matrices $\mathbf{Q}_{T}^{(i)}=\textbf{P}_{T}^{(i)}\textbf{W}^{Q(i)}_{T}$, key matrices $\mathbf{K}_{T}^{(i)}=\mathbb{W}_{trend}\textbf{W}^{K(i)}_{T}$, value matrices $\mathbf{V}_{T}^{(i)}=\mathbb{W}_{trend}\textbf{W}^{V(i)}_{T}$ for trend; query matrices $\mathbf{Q}_{S}^{(i)}=\textbf{P}_{S}\textbf{W}^{Q(i)}_{S}$, key matrices $\mathbf{K}_{S}^{(i)}=\textbf{E}^{{}^{\prime}}_{seasonal}\textbf{W}^{K(i)}_{S}$, value matrices $\mathbf{V}_{S}^{(i)}=\textbf{E}^{{}^{\prime}}_{seasonal}\textbf{W}^{V(i)}_{S}$ for seasonal; query matrices $\mathbf{Q}_{R}^{(i)}=\textbf{P}_{R}\textbf{W}^{Q(i)}_{R}$, key matrices $\mathbf{K}_{R}^{(i)}=\textbf{E}^{{}^{\prime}}_{residual}\textbf{W}^{K(i)}_{R}$, value matrices $\mathbf{V}_{R}^{(i)}=\textbf{E}^{{}^{\prime}}_{residual}\textbf{W}^{V(i)}_{R}$ for residual. Through multi-head attention, we reprogram each time series component.
$\mathbf{Z}_{T}^{(i)}=\text{ATTENTION}(\mathbf{Q}_{T}^{(i)},\mathbf{K}_{T}^{(i)},\mathbf{V}_{T}^{(i)})=\text{SOFTMAX}\left(\frac{\mathbf{Q}_{T}^{(i)}\left(\mathbf{K}_{T}^{(i)}\right)^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{T}^{(i)}$ (4)
where $d_{k}$ is the dimension of each head in the multi-head attention module.
3.4 Component-specific prompts & output projection
Prompting techniques effectively encode task-specific knowledge, improving model accuracy and coherence. However, translating time series into natural language is challenging, complicating instruction-following datasets and real-time prompting [13]. Recent advances show that prompts can enrich input context and guide reprogrammed time series patches [6]. To leverage semantic information, we propose a component-specific prefix prompting strategy, incorporating dataset context, statistical features, and task instructions for trend, seasonal, and residual components. For example, the instruction ’forecast the next 96 steps given the previous 512 steps [trend, seasonal, residual]’ serves as a template.
After processing these prompts through the backbone LLM, we retain the component embeddings and apply a linear projection. By denormalizing and summing these representations, we derive the final forecasts $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$.
@@ 4 Experiments
In our experiments, the proposed method outperforms state-of-the-art forecasting approaches across various benchmarks, including long-term, short-term, and few-shot forecasting. For a fair comparison, we follow the configurations outlined in [12] across all baselines, utilizing a unified evaluation pipeline [^1] https://github.com/thuml/Time-Series-Library.
Baselines. We compare with the SOTA time series models and cite their performance from [16, 3] if applicable. The baselines include Transformer-based models like PatchTST [7], non-transformer-based techniques, i.e., DLinear [15], and four LLM-based approaches: TimeLLM [6], LLM4TS [3], GPT4TS [16], and LLMTime [5]. Aligned with the GPT4TS configuration [16], we utilize only the first 6 layers of the 12-layer GPT-2 base [9] as the backbone for our model and TimeLLM.
Long-term Forecasting. For long-term forecasting, we evaluate on ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity(ECL), and Traffic, which have been widely adopted as benchmarking datasets for long-term forecasting works [12]. The input time series length $L$ is set as 512, and we evaluate across four prediction horizons: $H\in\{96,192,336,720\}$. The evaluation metrics include mean square error (MSE) and mean absolute error (MAE).
Table 1 presents the performance of various time series forecasting models on MSE and MAE metrics across different prediction horizons on multiple benchmarks. Our proposed model consistently outperforms existing baselines, demonstrating superior performance on average across most datasets and prediction lengths. This highlights the applicability of multi-level text alignment. Notably, our comparison with TimeLLM—a recent work leveraging text prototype reprogramming to align time series with text tokens—is significant. Specifically, our model achieves substantial improvements on the Weather and ETTm1 datasets, exceeding the best-performing model, LLM4TS, by 23.3% and 26.8%, respectively, in terms of MSE. Additionally, it records the lowest error rates across numerous dataset-prediction length configurations. These results suggest that integrating LLMs with multi-level text alignment can significantly enhance the accuracy of long-term time series forecasting. We also conduct a few-shot forecasting. More detailed experimental results for long-term forecasting and few-shot forecasting are in the appendix and full paper [1].
Table 1: Long-term forecasting results. All results are averaged from four different forecasting horizons: $H\in\{96,192,336,720\}$ horizons. Lower values indicate better performance. bold: the best, underline: second best.